Vasudevan, Inc. forecasts the free cash flows (in millions) shown below. If the weighted average cost of capital is 16% and the free cash flows are expected to continue growing at the same rate after Year 3 as from Year 2 to Year 3, what is the Year 0 value of operations, in millions? Enter your answer rounded to two decimal places. Do not enter $ or comma in the answer box. For example, if your answer is $12,300.456 then enter as 12300.46 in the answer box.                           

Year Free Cash Flow
1 $         (22.00)
2 $          42.00
3 $          45.00
Answer #1

Growth rate =(Free cash Flow in Year 3 - Free cash flow In Year 2)/Free cash Flow in Year 2 =(45-42)/42 = 7.1429%

Free Cash flow in Year 4 = Free cash Flow in Year 3*(1+Growth rate) =45*(1+7.1429%) =48.2143

Terminal value = Free cash Flow in Year 4/(Cost of Capital - Growth) =48.2143/(16%-7.1429%) =544.3548

PV of cash Flows = -22.00/(1+16%) +42/(1+16%)^2 +45/(1+16%)^3+544.3548/(1+16%)^3 =389.82